Sunday, July 02, 2017

fixed dew heating system

I fixed the Kendrick dew removal system in the Geoff Brown Observatory. It had not been working properly last night.

When I read the configuration from the hand controller to the computer, I found the same condition as last night. Lots of errors: AirT, Humi, and DewP. I ignored at the Opt2 error when Dietmar reminded me we only have the one objective sensor (on the C14 dew wrap). I noted there was still no power being directed to output 1 and 2.

I tested the dew heater wraps for the C14 and TV101 on channel 3 and was happy to find them working fine. Lots of heat.

I reseated the C14 wrap temperature sensor plug in the controller but it did not resolve any of the errors.

As I went to reseat the short temperature/humidity sensor on the controller, I found it loose. I fully seated it and the AirT, Humi, and DewP errors cleared. The controller started outputting to channel 1 and 2. Problem solved, at last.


The whole C14 dew wrap mounting was screwy so I redid it. I put the strap behind the end ring. I ensured the temperature sensor was under the strap.

I found the Tele Vue 101 dew heater around the University Optics finder scope. Moved it back to the proper refractor.

I installed a new elastic band from the kitchen and attached the flat finder heater to Telrad body. For many months, this had been flopping around on the top of the 'scope.

I remounted and secured the controller at the end of the Versa plate (it had been pulled out of position at some point).

I re-routed the eyepiece dew wrap wires to give us more slack.


As I had the Kendrick software running, I changed the controller from the winter to the summer profile (which is saved under configuration # 1). This uses the following duty cycles (the ON-time percentage):

C14 90%
TV101 35%
eyepiece A 75%
eyepiece B 75%

The C14 wrap sensor was configured to regulate the objective at 3 degrees above the air temperature. The TV101 channel was slaved to the C14.


Perhaps the temperature and humidity sensor had been bumped at some point causing it to be partially dislodged. However, it looked like the entire controller had been pulled out of its moorings. Was this due to a human pulling on cables or a cable snagging during a slew. Chris had noticed something odd but it had not been severe enough for his to report a problem. I was relieved nothing was broken. Reported my findings and outcome to the supervisors.

No comments: